OpenStudy (goblingnaw89):

x2 − 2x = 0 solve for x?

OpenStudy (dallasrichards16):

smile more bruhbruh

OpenStudy (wolf1728):

Using quadratic equation a=1 b=-2 c=0 x = [-b +- sqroot(b^2 - 4ac)] / 2a x = [--2 +-sqroot(4 - 4*1*0)] / 2*1 x = [2 +- sqroot(4) ]/ 2 That should help you

OpenStudy (girlie222):

2,0

OpenStudy (itsonlycdeee):

Don't need the quadratic equation. x^2 - 2x = 0 can be factored into x ( x - 2 ) = 0 . Hopefully that helps explain how the person above got 2 & 0.